OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (jim766):

find slope x1 x2 y1 y2 (3, 7) (-2, -8) \[m=\frac{ y2-y2 }{ x2-x1 }\]

OpenStudy (jim766):

parallel lines have the same slope....so this is the first step

OpenStudy (anonymous):

-8-7/-2-3?

OpenStudy (jim766):

right....keep going

OpenStudy (anonymous):

-15/-5

OpenStudy (anonymous):

would it be 3/1?

OpenStudy (jim766):

perfect...and like I said parallel lines have the same slope, so now you know that the slope of your parallel line is 3/1 or 3

OpenStudy (anonymous):

its not -3 right?

OpenStudy (jim766):

no it is 3, neg div neg is positive now the other thing you need for your parallel line is a y intercept instead of giving you that, the gave you a point. with that point we can find the y intercept y = mx + b which is the y intercept

OpenStudy (anonymous):

wait so 3 isnt the slope?

OpenStudy (jim766):

yes the slope, m = 3 x y (-2, 3) y = mx + b

OpenStudy (anonymous):

2=m(-2)+b?

OpenStudy (jim766):

close, remember we know the slope too we are going to find the yintercept, the b, by substituting in what we do know y = mx + b 3 = 3(-2) + b

OpenStudy (jim766):

now solve for b

OpenStudy (anonymous):

oh um, 9=b

OpenStudy (jim766):

right, now just put your slope m = 3 and the y intercept b = 9 together for your equation y = mx + b m = 3, b = 9

OpenStudy (jim766):

you good?